### v1

# Part I: General framework, application to Hartree-Fock and DFT

Abstract
Computing excited states of many-body quantum Hamiltonians is a fundamental challenge in computational physics and chemistry, with state-of-the-art methods broadly classified into variational (critical point search) and linear response approaches. This article introduces a unified framework based on the Kähler manifold formalism, which naturally accommodates both strategies for a wide range of variational models, including Hartree-Fock, CASSCF, Full CI, and adiabatic TDDFT. In particular, this formalism leads to a systematic and straightforward way to obtain the final equations of linear response theory for nonlinear models, which provides, in the case of mean-field models (Hartree-Fock and DFT), a simple alternative to Casida’s derivation. We detail the mathematical structure of Hamiltonian dynamics on Kähler manifolds, establish connections to standard quantum chemistry equations, and provide theoretical and numerical comparisons of excitation energy computation schemes at the Hartree-Fock level.
1 Introduction
Computing excited states of many-body quantum Hamiltonians is one of the biggest challenges in computational physics and chemistry 1.
The main approaches for computing approximate excited states can be classified into two categories:
  1. 1.
    variational or bivariational methods: computation of critical points of energy functionals on differentiable manifolds or algebraic varieties. This category encompasses state-specific and state-averaged methods 2, 3, 4, 5, 6, 7, 8;
  2. 2.
    methods based on linear response theory. Linear-response Hartree-Fock 9, 10, 11, 12, 13, 14, 15(LR-HF), Complete Active Space Self-Consistent Field 16, 17, 18, 19 (LR-CASSCF), time-dependent density functional theory 20, 21, 22, 23, 24, 25 (LR-TDDFT), and equation-of-motion coupled cluster 26, 27, 28, 29, 30, 31, 32, 33, 34 (EOM-CC) fall into this category.
All these approaches coincide for the reference electronic Schrödinger equation set on the continuous position space $\bigwedge^{N}L^{2}({\mathbb{R}}^{3};{\mathbb{C}}^{2})$, as well as for the Full Configuration Interaction (FCI) approximation 35, 36, but differ when further approximations are made. This article aims at providing a unified and compact framework covering critical point search (CP) and linear response theory (LR) from a theoretical viewpoint.
We will use the Kähler manifold formalism allowing one to conveniently formulate CP and LR for any variational model, including any variant of HF (restricted, unrestricted, general, restricted open-shell), CASSCF, and FCI methods, in a unified and compact framework. It also applies to adiabatic TDDFT 37, 38, as this method has an intrinsic Hamiltonian Kähler structure, although it is not a variational approximation of the reference electronic Schödinger equation. We will concentrate on the FCI and HF theories in this study. The cases of more advanced variational approximations such as CASSCF 39 and bivariational approximations such as CC 40, 41, 42 will be dealt with in a future work.
If $\mathcal{M}$ is a Kähler manifold (the definition and properties of this mathematical structure will be recalled in Section 2), and $\mathcal{E}:\mathcal{M}\to{\mathbb{R}}$ an energy functional over $\mathcal{M}$, the Hamiltonian Kähler dynamics is defined as
${\frac{dx}{dt}(t)=J_{x(t)}{\rm grad}_{\mathcal{M}}\mathcal{E}(x(t))},$ (1)
where ${\rm grad}_{\mathcal{M}}\mathcal{E}(x)$ is the Riemannian gradient of $\mathcal{E}$ at point $x\in\mathcal{M}$, and $J_{x}$ the so-called symplectic operator at point $x$. The critical points (steady states) therefore satisfy
${{\rm grad}_{\mathcal{M}}\mathcal{E}(x_{\star})=0}.$ (2)
The CP approach for computing excitation energies consists in:
  1. 1.
    finding critical points $x_{0},x_{1},x_{2},\ldots$ of $\mathcal{E}$ on $\mathcal{M}$ that can be identified with approximate ground and excited states, labelled in such a way that $\mathcal{E}(x_{0})\leq\mathcal{E}(x_{1})\leq\mathcal{E}(x_{2})\leq\cdots$;
  2. 2.
    computing the $k$-th excitation energy as the difference
    $\omega^{\rm CP}_{k}:=\mathcal{E}(x_{k})-\mathcal{E}(x_{0}),$ (3)
    between the energies of the approximation $x_{k}$ of the $k$-th excited state and the approximation $x_{0}$ of the ground state, i.e., the (presumably) global minimum of $\mathcal{E}$ on $\mathcal{M}$.
Step 1 is quite involved. Indeed, in contrast with linear models such as FCI, not all critical points of nonlinear models such as HF or CASSCF are suitable approximations of ground or excited states of the reference electronic Schrödinger equation. In general, the energy functional $\mathcal{E}$ can have many spurious critical points on $\mathcal{M}$. Worse still, there can be mismatch between the $k$-th excited states and the critical points with a Morse index $k$, i.e., critical points where the Riemannian Hessian has exactly $k$ negative eigenvalues. This phenomenon is illustrated on a 2D toy model in Fig. 1, as well as on an H4 molecule in Section 5.2. In Ref. 43, Lewin introduced a nonlinear min-max characterization of physically-relevant critical points for CASSCF with no internal orbitals. In Ref. 44, an algorithm for computing CASSCF approximations of first excited states based on this characterization was proposed and tested on an H2 molecule.
Figure 1: Contour of a 2D function $E:(0,1)\times(0,1)\to{\mathbb{R}}$ with one global minimizer ($A$), one local, non-global minimizer ($A^{\prime}$) and one saddle point with Morse index 1 ($B$). The restriction $\mathcal{E}$ of this function to the one-dimensional manifold $\mathcal{M}$ (indicated by the orange curve) has 3 global minimizers ($a_{1}$, $a_{2}$, $a^{\prime}$), 0 local, non-global, minimizers, and 2 critical points ($b$, $b^{\prime}$) with a Morse index 1, which are also local maxima since $\mathcal{M}$ is one-dimensional. The points $a_{1}$ and $a_{2}$ can be considered as reasonable approximations of $A$, $a^{\prime}$ as a reasonable approximation of $A^{\prime}$, $b$ as a reasonable approximation of $B$. The point $b^{\prime}$ is a spurious critical point, which does not approximate well any saddle point of $E$.
In LR, the excitation energies are identified with the resonant energies of the linearized dynamics around the global minimum $x_{0}$ of $\mathcal{E}$ on $\mathcal{M}$. This dynamics is given by
${\frac{dy}{dt}(t)=J_{x_{0}}{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{0})y(t)},$ (4)
where ${\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{0})$ is the Riemannian Hessian of $\mathcal{E}$ at $x_{0}$. Under a nondegeneracy assumption, it follows that the approximate excitation energies $\omega^{\rm LR}_{k}$ computed by LR are in fact the symplectic eigenvalues of ${\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{0})$:
$\{\omega_{k}^{\rm LR}\}_{1\leq k\leq\dim_{{\mathbb{C}}}(\mathcal{M})}:=\{\mbox{symplectic eigenvalues of }{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{0})\}.$ (5)
This approach provides a systematic and conceptually simple way to derive linear-response approximations to excitation energies for any variational model with Kähler structure.
In Section 2, we present the abstract framework of Hamiltonian dynamics on a generic Kähler manifold, and define precisely the mathematical objects arising in the fundamental Eqs. (1)-(5). In Section 3, we consider the specific case of Hamiltonian dynamics on Grassmann manifolds
$\mathcal{M}={\rm Gr}_{\mathbb{C}}(r,n):=\left\{\gamma\in{\mathbb{C}}^{n\times n}_{\rm herm}\;|\;\gamma^{2}=\gamma,\;{\rm Tr\,}(\gamma)=r\right\},$ (6)
allowing one to deal with FCI (or more generally any linear Schrödinger or Liouville equation on a finite-dimensional state-space), HF, and adiabatic TDDFT. In Section 3.1, we describe the Kähler structure of Grassmann manifolds. In Section 3.2, we give more explicit formulations of Eqs. (1)-(5) in the special case when $\mathcal{M}$ is a Grassmann manifold and apply this formalism to a molecular system with $N_{\rm b}$ atomic spin-orbitals and $N$ electrons. In Section 3.3, we deal with the FCI framework, where $n=\binom{N_{\rm b}}{N}$, $r=1$, $\Gamma\in{\rm Gr}_{\mathbb{C}}(1,n)\subset{\mathbb{C}}_{\rm herm}^{n\times n}$ represents the $N$-body density matrix, and the energy functional ${\mathcal{E}}$ is linear, i.e., ${\mathcal{E}}(\Gamma)={\rm Tr\,}(H\Gamma)$ with $H\in{\mathbb{C}}_{\rm herm}^{n\times n}$ the Hamiltonian matrix (specified later). We show that Eq. (1) is then equivalent to the time-dependent Schrödinger equation:
$\widehat{\Gamma}(t)=|\Psi(t)\rangle\langle\Psi(t)|\quad\mbox{with}\quad i\frac{d\Psi}{dt}(t)=\widehat{H}\Psi(t),\quad\|\Psi(t)\|=1,$
and that Eq. (2) is equivalent to the time-independent Schrödinger equation:
$\widehat{\Gamma}=|\Psi\rangle\langle\Psi|\quad\mbox{with}\quad\widehat{H}\Psi=E\Psi,\quad\|\Psi\|=1.$
We also show that, denoting by $E_{0}$ the ground-state energy of $\widehat{H}$, by $\Psi_{\star}$ a normalized ground state, and by $\widehat{\Gamma}_{\star}=\ket{\Psi_{\star}}\bra{\Psi_{\star}}$ |, the symplectic eigenvalues of ${\rm Hess}_{\mathcal{M}}\mathcal{E}(\Gamma_{\star})$ coincide with the eigenvalues of $(\widehat{H}-E_{0})|_{\Psi_{\star}^{\perp}}.$ It follows that, as expected,
$\omega_{k}^{\rm CP-FCI}=\omega_{k}^{\rm LR-FCI}=E_{k}-E_{0},$
where $E_{k}$ is the energy of the $k$-th FCI excited state. In Section 3.4, we consider the case of mean-field models, including HF and adiabatic TDDFT. In this case, $n=N_{\rm b}$, $r=N$, $\gamma\in{\mathbb{C}}_{\rm herm}^{N_{\rm b}\times N_{\rm b}}$ represents the one-body reduced density matrix (1-RDM), and the energy functional ${\mathcal{E}}$ is nonlinear. Eq. (1) is then equivalent to the time-dependent HF (TDHF) or adiabatic TDDFT equations, Eq. (2) to the usual HF or Kohn-Sham equations, and Eq. (5) leads to a straightforward derivation of the final equation providing single excitation energies. Recall that this equation is usually inferred from Casida’s equations of LR-TDHF or LR-TDDFT 23. In Section 4, we make use of the compact formalisms (1)-(5) to provide a theoretical comparison of the electronic excitation energy computation methods based on CP on the one-hand and LR on the other hand, at the UHF level of theory in the weakly interacting regime. More precisely, we consider a family of energy functionals of the form
$\mathcal{E}(\lambda,\gamma):={\rm Tr}(h\gamma)+\frac{\lambda}{2}{\rm Tr}(G(\gamma)\gamma)\quad\text{on}\quad\mathcal{M}={\rm Gr}_{{\mathbb{C}}}(N,N_{\rm b}),$
where $h\in{\mathbb{C}}^{N_{\rm b}\times N_{\rm b}}_{\rm herm}$ and $G:{\mathbb{C}}_{\rm herm}^{N_{\rm b}\times N_{\rm b}}\to{\mathbb{C}}_{\rm herm}^{N_{\rm b}\times N_{\rm b}}$ combines the Coulomb and exchange terms (defined later). We compare the UHF electronic excitation energies $\omega^{\rm HF,CP}_{k}(\lambda)$ and $\omega^{\rm HF,LR}_{k}(\lambda)$ obtained by CP and LR respectively for small values of $\lambda$. In Section 5, we report numerical results supporting our theoretical analysis.
2 Hamiltonian dynamics on Kähler manifolds
In this section, we intend to provide a brief introduction to Kähler manifolds and associated Hamiltonian dynamics. We refer interested readers to e.g., the monographs  45, 46, 47.
A Kähler manifold is a complex manifold $\mathcal{M}$ of complex dimension $n={\rm dim}_{\mathbb{C}}(\mathcal{M})$, endowed with a positive definite Hermitian form $\langle\bullet,\bullet\rangle_{\bullet}$, which allows one to endow $\mathcal{M}_{\mathbb{R}}$ (the manifold $\mathcal{M}$ seen this time as a real $2n$-dimensional manifold) with
  1. 1.
    a Riemannian structure with metric defined by
    $\forall u,v\in T_{x}\mathcal{M}_{\mathbb{R}},\quad g_{x}(u,v):=\mbox{Re}(\langle u,v\rangle_{x});$
  2. 2.
    a symplectic structure with symplectic form defined by
    $\forall u,v\in T_{x}\mathcal{M}_{\mathbb{R}},\quad\omega_{x}(u,v):=\mbox{Im}(\langle u,v\rangle_{x}).$
The symplectic operator $J_{x}:T_{x}\mathcal{M}_{\mathbb{R}}\to T_{x}\mathcal{M}_{\mathbb{R}}$ is defined by
$\forall u,v\in T_{x}\mathcal{M}_{\mathbb{R}},\quad g_{x}(u,v)=\omega_{x}(u,J_{x}v).$
It is easy to check that, for any $u$, $v\in T_{x}\mathcal{M}_{{\mathbb{R}}}$,
$g_{x}(J_{x}u,J_{x}v)=g_{x}(u,v),\quad\omega_{x}(J_{x}u,J_{x}v)=\omega_{x}(u,v),\quad\omega_{x}(u,v)=-g_{x}(u,J_{x}v).$ (7)
Let $\mathcal{E}:\mathcal{M}_{\mathbb{R}}\to{\mathbb{R}}$ be a smooth real-valued function on $\mathcal{M}_{\mathbb{R}}$. We denote respectively by
$d\mathcal{E}:\mathcal{M}_{\mathbb{R}}\to T^{*}\mathcal{M}_{\mathbb{R}}\qquad\mbox{and}\qquad{\rm grad}_{\mathcal{M}}\mathcal{E}:\mathcal{M}_{\mathbb{R}}\to T\mathcal{M}_{\mathbb{R}}$
the differential of $\mathcal{E}$ and its Riemannian gradient. For each $x\in\mathcal{M}_{\mathbb{R}}$, $d\mathcal{E}(x)$ is the unique ${\mathbb{R}}$-linear form on $T_{x}\mathcal{M}_{\mathbb{R}}$ such that for any smooth path $c:[-1,1]\to\mathcal{M}_{\mathbb{R}}$ drawn on $\mathcal{M}_{\mathbb{R}}$ satisfying $c(0)=x$ and $\dot{c}(0)=v\in T_{x}\mathcal{M}_{\mathbb{R}}$,
$\mathcal{E}(c(t))=\mathcal{E}(x)+t\,d\mathcal{E}(x)(v)+o(t),$
and ${\rm grad}_{\mathcal{M}}E(x)$ the unique vector of $T_{x}\mathcal{M}_{\mathbb{R}}$ such that
$\forall v\in T_{x}\mathcal{M}_{\mathbb{R}},\quad d\mathcal{E}(x)(v)=g_{x}\left({\rm grad}_{\mathcal{M}}\mathcal{E}(x),v\right).$
The Hamiltonian dynamics on the symplectic manifold $(\mathcal{M}_{\mathbb{R}},J_{\bullet})$ is defined by
${\frac{dx}{dt}(t)=J_{x(t)}{\rm grad}_{\mathcal{M}}\mathcal{E}(x(t))}.$ (8)
The equilibrium points (steady states) satisfy
${{\rm grad}_{\mathcal{M}}\mathcal{E}(x_{\star})=0}$ (9)
and the LR around an equilibrium point $x_{\star}$ consists in studying the linearized dynamics around $x_{\star}$ given by
${\frac{dy}{dt}(t)=J_{x_{\star}}{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})y(t)},$ (10)
where ${\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star}):T_{x_{\star}}\mathcal{M}_{\mathbb{R}}\to T_{x_{\star}}\mathcal{M}_{\mathbb{R}}$ is the Riemannian Hessian of $\mathcal{E}$ at $x_{\star}$. Recall that ${\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})$ is the unique symmetric ${\mathbb{R}}$-linear operator on $T_{x_{\star}}\mathcal{M}_{\mathbb{R}}$ such that for all smooth paths $c:[-1,1]\to\mathcal{M}_{\mathbb{R}}$ satisfying $c(0)=x_{\star}$ and $\dot{c}(0)=v\in T_{x_{\star}}\mathcal{M}_{\mathbb{R}}$,
$\mathcal{E}(c(t))=\mathcal{E}(x_{\star})+\frac{t^{2}}{2}\,g_{x_{\star}}(v,{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})v)+o(t^{2}),$
and that symmetric means
$\forall u,v\in T_{x_{\star}}\mathcal{M}_{\mathbb{R}},\quad g_{x_{\star}}(u,{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})v)=g_{x_{\star}}({\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})u,v).$
Eq. (10) is a Hamiltonian dynamics on the symplectic space $(T_{x_{\star}}\mathcal{M}_{\mathbb{R}},J_{x_{\star}})$ with quadratic Hamiltonian
$\mathfrak{h}_{x_{\star}}(y):=\frac{1}{2}g_{x_{\star}}(y,{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})y).$
Let $(u_{1},\ldots,u_{n})$ be an orthonormal basis of $T_{x_{\star}}\mathcal{M}$ for $\langle\bullet,\bullet\rangle_{x_{\star}}$, i.e., a family of $n$ vectors of $T_{x_{\star}}\mathcal{M}$ such that
$\langle u_{j},u_{j^{\prime}}\rangle_{x_{\star}}=\delta_{jj^{\prime}},$
and $\mathcal{X}:={\rm Span}_{\mathbb{R}}(u_{1},\ldots,u_{n})$ the $n$-dimensional real-vector space spanned by the $u_{j}$’s. We have
$T_{x_{\star}}\mathcal{M}_{\mathbb{R}}=\mathcal{X}\oplus_{\mathbb{R}}(J_{x_{\star}}\mathcal{X}),$
and we infer from Eq. (7) that in the basis $(u_{1},\ldots,u_{n},{J_{x_{\star}}u_{1},\ldots,J_{x_{\star}}u_{n}})$ of $T_{x_{\star}}\mathcal{M}_{\mathbb{R}}$, $g_{x_{\star}}$ is represented by the identity matrix $I_{2n}$, $J_{x_{\star}}$ by the skew symmetric matrix
$J:=\begin{pmatrix}0&-I_{n}\\ I_{n}&0\end{pmatrix}.$
The Hessian matrix ${\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})$ is represented by the real symmetric matrix
$\mathfrak{H}_{\star}:=\left(\begin{array}[]{cc}H^{\rm rr}&H^{\rm ri}\\ H^{\rm ir}&H^{\rm ii}\end{array}\right)\in{\mathbb{R}}^{2n\times 2n}_{\rm sym}$ (11)
with entries
$[H^{\rm rr}]_{jj^{\prime}}$ $:=g_{x_{\star}}(u_{j},{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})u_{j^{\prime}}),$ (12)
$[H^{\rm ri}]_{jj^{\prime}}$ $:=g_{x_{\star}}(u_{j},{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})(J_{x_{\star}}u_{j^{\prime}})),$ (13)
$[H^{\rm ir}]_{jj^{\prime}}$ $:=g_{x_{\star}}(J_{x_{\star}}u_{j},{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})u_{j^{\prime}}),$ (14)
$[H^{\rm ii}]_{jj^{\prime}}$ $:=g_{x_{\star}}(J_{x_{\star}}u_{j},{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})(J_{x_{\star}}u_{j^{\prime}})).$ (15)
If $x_{\star}$ is a nondegenerate local minimizer of $\mathcal{E}$ on $\mathcal{M}_{\mathbb{R}}$, then the matrix $\mathfrak{H}_{\star}$ is real symmetric positive definite (SPD). According to Williamson’s theorem 48, there exists a symplectic matrix $S\in{\mathbb{R}}^{2n\times 2n}$ (i.e., a matrix in ${\mathbb{R}}^{2n\times 2n}$ such that $S^{T}JS=J$) such that
$S^{T}\mathfrak{H}_{\star}S=\left(\begin{array}[]{cc}\Omega&0\\ 0&\Omega\end{array}\right)\quad\mbox{with}\quad\Omega:=\mbox{diag}(\omega_{1}^{\rm LR},\ldots,\omega_{n}^{\rm LR}),\quad 0<\omega_{1}^{\rm LR}\leq\cdots\leq\omega_{n}^{\rm LR}.$ (16)
The $\omega_{j}^{\rm LR}$’s are called the symplectic eigenvalues of $\mathfrak{H}_{\star}$. It is easy to show that these quantities are independent of the chosen orthonormal basis $(u_{1},\ldots,u_{n})$ of $T_{x_{\star}}\mathcal{M}$. They are therefore intrinsic and are called the symplectic eigenvalues of ${\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})$.
Denoting by $Y(t)\in{\mathbb{R}}^{2n}$ the column vector containing the coefficients of $y(t)$ in the basis $(u_{1},\ldots,u_{n},J_{x_{\star}}u_{1},\ldots,J_{x_{\star}}u_{n})$ and by
$Z(t):=\left(\begin{array}[]{cc}\Omega^{-1/2}&0\\ 0&\Omega^{1/2}\end{array}\right)S^{-1}Y(t)=:\left(\begin{array}[]{c}q(t)\\ p(t)\end{array}\right)\quad\mbox{with}\quad q(t),p(t)\in{\mathbb{R}}^{n},$
we have
$\frac{dZ}{dt}(t)=J\nabla\mathfrak{H}^{\rm HO}_{\star}(Z(t))\quad\mbox{with}\quad\mathfrak{H}^{\rm HO}_{\star}(Z)=\mathfrak{H}^{\rm HO}_{\star}\left(\begin{array}[]{c}q\\ p\end{array}\right)=\frac{1}{2}|p|^{2}+\frac{1}{2}q^{T}\Omega^{2}q.$
The angular frequencies $\omega_{1},\ldots,\omega_{n}$ therefore represent the excitation frequencies of the system in the linear response regime.
Let us finally consider the special case of separable Hamiltonians of the form
$\mathfrak{H}_{\star}=\left(\begin{array}[]{cc}H^{\rm rr}&0\\ 0&H^{\rm ii}\end{array}\right)\in{\mathbb{R}}^{2n\times 2n}_{\rm sym}.$ (17)
Introducing the symplectic matrix
$S=\left(\begin{array}[]{cc}R&0\\ 0&R^{-T}\end{array}\right)\quad\mbox{with}\quad R=(H^{\rm rr})^{-1/2}\left((H^{\rm rr})^{1/2}H^{\rm ii}(H^{\rm rr})^{1/2}\right)^{1/4},$
we get
$S^{T}\mathfrak{H}_{\star}S=\left(\begin{array}[]{cc}\widetilde{\Omega}&0\\ 0&\widetilde{\Omega}\end{array}\right),\quad\mbox{with}\quad\widetilde{\Omega}=\left((H^{\rm rr})^{1/2}H^{\rm ii}(H^{\rm rr})^{1/2}\right)^{1/2}.$ (18)
Diagonalizing $\widetilde{\Omega}$ readily provides the symplectic eigenvalues of $\mathfrak{H}_{\star}$.
3 Dynamics on Grassmann manifolds (FCI, TDHF, TDDFT)
In this section, we restrict ourselves to the case when the Kähler manifold is the complex Grassmann manifold
$\mathcal{M}={\rm Gr}_{\mathbb{C}}(r,n)=\left\{\gamma\in{\mathbb{C}}^{n\times n}_{\rm herm}\;|\;\gamma^{2}=\gamma,\;{\rm Tr\,}(\gamma)=r\right\},$ (19)
where $r,n$ are positive integers such that $11\leq r\leq n-$ 1 (the trivial cases $\mathcal{M}={\rm Gr}_{\mathbb{C}}(0,n)=\{0\}$ and $\mathcal{M}={\rm Gr}_{\mathbb{C}}(n,n)=\{I_{n}\}$ are uninteresting). Such manifolds appear in quantum dynamics in at least two settings:
  1. 1.
    pure-state dynamics of generic isolated finite-dimensional quantum systems described by the linear Schrödinger equation, or equivalently the linear quantum Liouville equation (see Section 3.3). In computational quantum chemistry, this corresponds to the FCI level of theory;
  2. 2.
    mean-field models for many-body fermionic systems. This framework encompasses the cases of both TDHF and adiabatic TDDFT, for which the dynamics is governed by a system of coupled nonlinear Schrödinger equations, or equivalently a nonlinear quantum Liouville equation (see Section 3.4).
Before elaborating on these two fundamental examples, the geometry and Kähler structure of the complex Grassmann manifold ${\rm Gr}_{\mathbb{C}}(r,n)$ are recalled in Sections 3.1 and 3.2, and the abstract Eqs. (8)-(10) are rewritten in a more concrete form in the special case when $\mathcal{M}={\rm Gr}_{\mathbb{C}}(r,n)$ in Sections 3.3 and 3.4. More details about Grassmann manifold can be found in the review articles in Refs. 49, 50 and the monographs in Refs. 51, 52.
3.1 Geometry of complex Grassmann manifolds
The tangent space to the complex Grassmann manifold $\mathcal{M}={\rm Gr}_{\mathbb{C}}(r,n)$ at some $\gamma\in\mathcal{M}$ is given by
$T_{\gamma}\mathcal{M}=\{Q\in{\mathbb{C}}^{n\times n}_{\rm herm}\;|\;\gamma Q+Q\gamma=Q\}.$
Any $\gamma\in\mathcal{M}$ can be decomposed as
$\gamma=U\left(\begin{array}[]{cc}I_{r}&0\\ 0&0\end{array}\right)U^{*},\quad\mbox{with}\quad UU^{*}=U^{*}U=I_{n},$
and we then have
$T_{\gamma}\mathcal{M}=\left\{\left.Q=U\left(\begin{array}[]{cc}0&X\\ X^{*}&0\end{array}\right)U^{*}\right|X\in{\mathbb{C}}^{r\times m}\right\}\quad\mbox{with}\quad m:=n-r.$ (20)
The Grassmann manifold $\mathcal{M}$ can be endowed with the following Kähler structure:
  • in view of Eq. (20), the tangent space $T_{\gamma}\mathcal{M}$ can be identified with the complex vector space ${\mathbb{C}}^{r\times m}$, endowed with the Hermitian inner product
    $\forall\gamma\in\mathcal{M},\quad\forall Q_{1},Q_{2}\in T_{\gamma}\mathcal{M},\quad\langle Q_{1},Q_{2}\rangle_{\gamma}:=2{\rm Tr\,}(X_{1}^{*}X_{2})=2{\rm Tr\,}\left(Q_{1}\gamma Q_{2}(1-\gamma)\right),$
    where $X_{j}$ is the matrix of ${\mathbb{C}}^{r\times m}$ such that
    $Q_{j}=U\left(\begin{array}[]{cc}0&X_{j}\\ X_{j}^{*}&0\end{array}\right)U^{*},\quad j=1,2;$ (21)
  • Riemannian metric: using the properties
    $\gamma^{2}=\gamma,\quad(1-\gamma)^{2}=1-\gamma,\quad\forall Q\in T_{\gamma}\mathcal{M}_{\mathbb{R}},\;\gamma Q+Q\gamma=Q\mbox{ i.e., }\gamma Q=Q(1-\gamma),$ (22)
    we get, for any $\gamma\in\mathcal{M}_{{\mathbb{R}}}$ and $Q_{1},Q_{2}\in T_{\gamma}\mathcal{M}_{{\mathbb{R}}}$,
    $g_{\gamma}(Q_{1},Q_{2}):$ $={\rm Re}(\langle Q_{1},Q_{2}\rangle_{\gamma})$
    $={\rm Tr\,}(X_{1}^{*}X_{2})+{\rm Tr\,}(X_{1}X_{2}^{*})$
    $={\rm Tr\,}\left(Q_{1}\gamma Q_{2}(1-\gamma)\right)+{\rm Tr\,}\left((1-\gamma)Q_{2}\gamma Q_{1}\right)$
    $={\rm Tr\,}\left(Q_{1}\gamma Q_{2}\right)+{\rm Tr\,}\left(Q_{1}(1-\gamma)Q_{2}\right)$
    $={\rm Tr\,}(Q_{1}Q_{2});$
  • Symplectic form: we infer from Eq. (22) that, for any $\gamma\in\mathcal{M}_{{\mathbb{R}}}$ and $Q_{1},Q_{2}\in T_{\gamma}\mathcal{M}_{{\mathbb{R}}}$,
    $\omega_{\gamma}(Q_{1},Q_{2}):$ $={\rm Im}(\langle Q_{1},Q_{2}\rangle_{\gamma})$
    $=-i{\rm Tr\,}\left(Q_{1}(2\gamma-1)Q_{2}\right)$
    $=-i{\rm Tr\,}\left(Q_{2}(1-2\gamma)Q_{1}\right).$
    Note that, still using Eq. (22), we have
    $\overline{{\rm Tr\,}(Q_{1}(2\gamma-1)Q_{2})}$ $={\rm Tr\,}(Q_{2}(2\gamma-1)Q_{1})=2{\rm Tr\,}(Q_{2}\gamma Q_{1})-{\rm Tr\,}(Q_{2}Q_{1})$
    $=2{\rm Tr\,}((1-\gamma)Q_{2}Q_{1}(1-\gamma))-{\rm Tr\,}(Q_{2}Q_{1})$
    $={\rm Tr\,}(Q_{1}(1-2\gamma)Q_{2})=-{\rm Tr\,}(Q_{1}(2\gamma-1)Q_{2}),$
    so that ${\rm Tr\,}(Q_{1}(2\gamma-1)Q_{2})$ is an imaginary number.
A simple calculation shows that the symplectic operator $J_{\gamma}:T_{\gamma}\mathcal{M}_{\mathbb{R}}\to T_{\gamma}\mathcal{M}_{\mathbb{R}}$ is given by
$\forall Q\in T_{\gamma}\mathcal{M}_{\mathbb{R}},\quad J_{\gamma}Q:=-i[Q,\gamma].$ (23)
In the representation (21), it reads
$\forall Q=U\left(\begin{array}[]{cc}0&X\\ X^{*}&0\end{array}\right)U^{*}\in T_{\gamma}\mathcal{M}_{\mathbb{R}},\qquad J_{\gamma}Q=U\left(\begin{array}[]{cc}0&iX\\ -iX^{*}&0\end{array}\right)U^{*}.$
Recall that the orthogonal projector $P_{\gamma}:{\mathbb{C}}^{n\times n}_{\rm herm}\to T_{\gamma}\mathcal{M}_{\mathbb{R}}$ (for the Frobenius inner product) has a simple expression:
$\forall\gamma\in\mathcal{M}_{\mathbb{R}},\quad\forall M\in{\mathbb{C}}^{n\times n}_{\rm herm},\quad P_{\gamma}M:=\gamma M(1-\gamma)+(1-\gamma)M\gamma=[\gamma,[\gamma,M]],$ (24)
and that in the representation (21),
$\forall M=U\left(\begin{array}[]{cc}A&B\\ B^{*}&C\end{array}\right)U^{*}\in{\mathbb{C}}^{n\times n}_{\rm herm},\quad P_{\gamma}M=U\left(\begin{array}[]{cc}0&B\\ B^{*}&0\end{array}\right)U^{*}.$
It follows by direct calculations that
$\forall\gamma\in\mathcal{M}_{\mathbb{R}},\quad\forall M\in{\mathbb{C}}^{n\times n}_{\rm herm},\quad$ $i[\gamma,M]\in T_{\gamma}\mathcal{M}_{\mathbb{R}},\quad[\gamma,[\gamma,[\gamma,M]]]=[\gamma,M].$ (25)
Finally, the exponential map is given by
$\forall\gamma\in\mathcal{M}_{\mathbb{R}},\quad\forall Q\in T_{\gamma}\mathcal{M}_{\mathbb{R}},\quad{\rm Exp}_{\gamma}(Q)=e^{[\gamma,Q]}\gamma e^{-[\gamma,Q]},$
which in essence coincides with the exponential parametrization of molecular coefficients widely used in quantum chemistry 53.
3.2 Hamiltonian dynamics on Grassmann manifolds
Let $\mathcal{M}={\rm Gr}_{\mathbb{C}}(r,n)$ be a complex Grassmann manifold, $E:{\mathbb{C}}^{n\times n}_{\rm herm}\to{\mathbb{R}}$ an energy functional, smooth in a neighborhood of $\mathcal{M}_{\mathbb{R}}$, and $\mathcal{E}:\mathcal{M}_{\mathbb{R}}\to{\mathbb{R}}$ the restriction of $E$ to $\mathcal{M}_{\mathbb{R}}$. Denoting by $H_{\gamma}:=\nabla E(\gamma)\in{\mathbb{C}}^{n\times n}_{\rm herm}$ the gradient of $E$ at $\gamma\in\mathcal{M}_{\mathbb{R}}$ (for the Frobenius inner product), we have
${\rm grad}_{\mathcal{M}}\mathcal{E}(\gamma)=P_{\gamma}(\nabla E(\gamma))=P_{\gamma}(H_{\gamma})=[\gamma,[\gamma,H_{\gamma}]].$ (26)
Using Eqs. (23) and (25), we get
$\forall\gamma\in\mathcal{M}_{\mathbb{R}},\quad J_{\gamma}{\rm grad}_{\mathcal{M}}\mathcal{E}(\gamma)=-i[{\rm grad}_{\mathcal{M}}\mathcal{E}(\gamma),\gamma]=-i[[\gamma,[\gamma,H_{\gamma}]],\gamma]=-i[H_{\gamma},\gamma].$
The symplectic dynamics (8) generated by the energy functional $\mathcal{E}$ is thus given by
[…]
Stationary points $\gamma_{\star}$ are characterized by the equation
[…]
This equilibrium condition implies that there exists a unitary matrix $U_{\star}\in U(n)$ such that
$\gamma_{\star}=U_{\star}\left(\begin{array}[]{cc}I_{r}&0\\ 0&0\end{array}\right)U_{\star}^{*}\quad\mbox{and}\quad H_{\gamma_{\star}}=U_{\star}\left(\begin{array}[]{cc}E^{r}&0\\ 0&E^{m}\end{array}\right)U_{\star}^{*},$ (29)
where $E^{r}\in{\mathbb{C}}^{r\times r}_{\rm herm}$ and $E^{m}\in{\mathbb{C}}^{m\times m}_{\rm herm}$ are real diagonal matrices.
As $\mathcal{M}$ is a Riemannian submanifold of ${\mathbb{C}}^{n\times n}_{\rm herm}$, the Riemannian Hessian on $\mathcal{E}$ is given by
$\forall Q\in T_{\gamma_{\star}}\mathcal{M}_{\mathbb{R}},\quad{\rm Hess}_{\mathcal{M}}\mathcal{E}(\gamma_{\star})(Q)=P_{\gamma_{\star}}([D\mathcal{G}(\gamma_{\star})](Q)),$
where $\mathcal{G}:{\mathbb{C}}^{n\times n}_{\rm herm}\to{\mathbb{C}}^{n\times n}_{\rm herm}$ is any smooth extension of ${\rm grad}_{\mathcal{M}}\mathcal{E}(\gamma)$ in a neighborhood of $\mathcal{M}_{\mathbb{R}}$. In view of (26), an obvious extension is
$\mathcal{G}(M):=[M,[M,\nabla E(M)]],$
for which
[…]
where $D^{2}E(\gamma_{\star}):{\mathbb{C}}^{n\times n}_{\rm herm}\to{\mathbb{C}}^{n\times n}_{\rm herm}$ is the second derivative of $E$ at $\gamma_{\star}$. The Riemannian Hessian ${\rm Hess}_{\mathcal{M}}\mathcal{E}(\gamma_{\star}):T_{\gamma_{\star}}\mathcal{M}_{\mathbb{R}}\to T_{\gamma_{\star}}\mathcal{M}_{\mathbb{R}}$ is thus given by
${\rm Hess}_{\mathcal{M}}\mathcal{E}(\gamma_{\star})(Q)=[\gamma_{\star},[Q,H_{\gamma_{\star}}]]+P_{\gamma_{\star}}\left(D^{2}E(\gamma_{\star})(Q)\right).$ (30)
Using Eqs. (25) and (28), we finally obtain that the linearized dynamics is given by
[…]
where we used the fact that Eqs. (22) and (23) imply
$\forall Q\in T_{\gamma_{\star}}\mathcal{M}_{\mathbb{R}},\quad J_{\gamma_{\star}}\left([\gamma_{\star},[Q,H_{\gamma_{\star}}]]\right)=-i[[\gamma_{\star},[Q,H_{\gamma_{\star}}]],\gamma_{\star}]=-i[H_{\gamma_{\star}},Q].$
3.3 Linear energy functionals on projective spaces (FCI)
Consider an isolated quantum system with finite-dimensional state space $\mathcal{H}$. In the quantum chemistry applications we have in mind, ${\mathcal{H}}$ is the FCI $N$-electron state-space generated by the $N_{\rm b}$ atomic spin-orbitals of the discretization basis, and is therefore of dimension $n=\binom{N_{\rm b}}{N}$.
The space $\mathcal{H}$ can be identified with ${\mathbb{C}}^{n}$ by the choice of an orthonormal basis. The postulates of quantum mechanics imply that the set of FCI pure states of the system can be identified with the projective space $P(\mathcal{H})\cong{\rm CP}^{n-1}\cong{\rm Gr}(1,n)$. Recall that an element of ${\rm CP}^{n-1}$ can be represented by
  1. 1.
    a normalized vector $\psi\in{\mathbb{C}}^{n}$, it being understood that $\psi$ and $e^{i\theta}\psi$, with any $\theta\in{\mathbb{R}}$, represent the same element of ${\rm CP}^{n-1}$. This representation is the usual one in quantum mechanics (wavefunction formalism);
  2. 2.
    a non-zero vector $\widetilde{\psi}\in{\mathbb{C}}^{n}$, it being understood that $\widetilde{\psi}$ and $\lambda\widetilde{\psi}$, with any $\lambda\in{\mathbb{C}}\setminus\{0\}$, represent the same element of ${\rm CP}^{n-1}$;
  3. 3.
    a unique element $\Gamma\in{\rm Gr}_{\mathbb{C}}(1,n)$. This representation is used in the density-matrix formulation of quantum mechanics.
The connection between these three representations is
$\Gamma=\psi\psi^{*}=\frac{\widetilde{\psi}\widetilde{\psi}^{*}}{|\widetilde{\psi}|^{2}}\quad\mbox{or, using Dirac's bra-ket notation,}\quad\Gamma=|\psi\rangle\langle\psi|=\frac{|\widetilde{\psi}\rangle\langle\widetilde{\psi}|}{\langle\widetilde{\psi}|\widetilde{\psi}\rangle}.$
Let $H\in{\mathbb{C}}^{n\times n}_{\rm herm}$ be the Hamiltonian matrix of the system in the chosen orthonormal basis of the FCI $N$-electron state-space. Let $E:{\mathbb{C}}^{n\times n}_{\rm herm}\to{\mathbb{R}}$ be the linear energy functional defined by $E(\Gamma)={\rm Tr\,}(H\Gamma)$. The FCI energy functional is the restriction $\mathcal{E}$ of $E$ to ${\rm Gr}_{\mathbb{C}}(1,n)$:
$\mathcal{E}(\Gamma)={\rm Tr\,}(H\Gamma)=\langle\psi|H|\psi\rangle=\frac{\langle\widetilde{\psi}|H|\widetilde{\psi}\rangle}{\langle\widetilde{\psi}|\widetilde{\psi}\rangle}.$
The FCI approximation is thus a special case of the general framework described in Section 3.2, corresponding to $r=1$, $n=\binom{N_{\rm b}}{N}$, and $E:{\mathbb{C}}^{n\times n}_{\rm herm}\to{\mathbb{R}}$ linear. The dynamics (27) then reads
$\frac{d\Gamma}{dt}(t)=-i[H,\Gamma(t)],\quad i\frac{d\psi}{dt}(t)=H\psi(t)-\beta(t)\psi(t),\quad i\frac{d\widetilde{\psi}}{dt}(t)=H\widetilde{\psi}(t)-\mu(t)\widetilde{\psi}(t),$ (32)
where $\beta(t)\in{\mathbb{R}}$, $\mu(t)\in{\mathbb{C}}$ are gauge parameters. Steady states (see Eq. (28)) are characterized by
$[H,\Gamma_{\star}]=0,\qquad\left\{\begin{array}[]{l}H\psi_{\star}=E\psi_{\star},\\ \|\psi_{\star}\|=1,\\ \psi_{\theta}(t)=e^{i\theta(t)}\psi_{\star},\end{array}\right.,\qquad\left\{\begin{array}[]{l}H\widetilde{\psi}_{\star}=E\widetilde{\psi}_{\star},\\ \widetilde{\psi}_{\lambda}(t)=e^{\lambda(t)}\psi_{\star}.\end{array}\right.$
In particular, the local minimizers of $\mathcal{E}$ are global and are the ground-states of $H$ and the assumption that the ground-state $\Gamma_{\star}$ is nondegenerate is equivalent to the assumption that the ground-state eigenvalue $E_{0}$ of $H$ is simple. Other critical points of $\mathcal{E}$ are then exactly the excited states of $H$, corresponding to the energies $E_{1}\leq E_{2}\leq\cdots\leq E_{n-1}$, and the excitation energies determined by CP are $\omega^{\rm CP-FCI}_{k}=E_{k}-E_{0}$ ($k=1,\ldots,n-1$).
Let us turn to LR. Due to the fact that the energy functional $E(\Gamma)={\rm Tr\,}(H\Gamma)$ is linear in $\Gamma$, we have $H_{\Gamma_{\star}}=\nabla E(\Gamma_{\star})=H$ and $D^{2}E(\Gamma_{\star})=0$. It follows that the linear response Eq. (31) around a steady state $\Gamma_{\star}=|\Psi_{\star}\rangle\langle\Psi_{\star}|$ on ${\rm Gr}_{\mathbb{C}}(1,n)$ (with $H\Psi_{\star}=E_{0}\Psi_{\star}$) reads
$\frac{dQ}{dt}(t)=-i[H,Q(t)]\quad\mbox{on}\quad T_{\Gamma_{\star}}{\rm Gr}_{\mathbb{C}}(1,n)_{\mathbb{R}}.$ (33)
Using an orthonormal basis of the FCI Hamiltonian, in which the first basis vector is $\Psi_{\star}$, to identify the discretized $N$-particle state space with ${\mathbb{C}}^{n}$, $n=\binom{N_{\rm b}}{N}$, the representation in Eq. (29) for $r=1$ can be rewritten as
$\Gamma_{\star}=\left(\begin{array}[]{cc}1&0\\ 0&0\end{array}\right),\quad Q=\left(\begin{array}[]{cc}0&\Psi^{*}\\ \Psi&0\end{array}\right),\quad\Psi\in{\mathbb{C}}^{n-1},$
$-i[H,Q]=\left(\begin{array}[]{cc}0&((H|_{\Psi_{\star}^{\perp}}-E_{0})\Psi)^{*}\\ (H|_{\Psi_{\star}^{\perp}}-E_{0})\Psi&0\end{array}\right),$
where $H|_{\Psi_{\star}^{\perp}}\in{\mathbb{C}}_{\rm herm}^{(n-1)\times(n-1)}$ denotes the projection of $H$ onto $\Psi_{\star}^{\perp}$. Consequently, Eq. (33) is equivalent to
$i\frac{d\Psi}{dt}(t)=(H|_{\Psi_{\star}^{\perp}}-E_{0})\Psi(t)\quad\mbox{on}\quad\Psi_{\star}^{\perp},$
which shows that the LR-FCI excitation energies are given as expected by $\omega^{\rm LR-FCI}_{k}=E_{k}-E_{0}$ ($k=1,\ldots,n-1$).
In summary, we recover the well-known fact that both CP and LR are exact for linear quantum Hamiltonians (FCI level of theory in quantum chemistry).
3.4 Mean-field models (TDHF and adiabatic TDDFT)
Let us now consider the case of mean-field models. We will focus on general HF (GHF) models, in which non-collinear spin states are allowed. However, the formalism presented here can be easily adapted to restricted and unrestricted HF models, more widely used in applications, as well as restricted/unrestricted adiabatic TDDFT.
We still denote by $N_{\rm b}$ the number of atomic spin-orbitals of the chosen discretization basis, and assume that these functions are real-valued, which is the case for standard atomic orbital basis sets. We also assume that the system is not subjected to an external magnetic field and that spin-orbit coupling terms are neglected. In this case, it makes sense to seek HF or Kohn-Sham ground states satisfying time-reversal symmetry, leading to real-valued HF or Kohn-Sham occupied orbitals, and real-symmetric Fock or Kohn-Sham matrices at the ground state. In the rest of the section, we work in an orthonormal basis of the one-particle state-space $\mathcal{H}$ consisting of $N_{\rm b}$ real occupied/virtual molecular spin-orbitals $(\varphi_{1},\ldots,\varphi_{N_{b}})$ at the GHF ground state $\gamma_{\star}$ with energies ranked in nondecreasing order. We denote by $h\in{\mathbb{C}}^{N_{\rm b}\times N_{\rm b}}_{\rm herm}$ the one-electron Hamiltonian matrix with
[…]
and by
$(pq|rs):=\int_{{\mathbb{R}}^{3}\times\{\alpha,\beta\}}\int_{{\mathbb{R}}^{3}\times\{\alpha,\beta\}}\frac{\overline{\varphi_{p}(\bm{x})}\varphi_{q}(\bm{x})\overline{\varphi_{r}(\bm{x}^{\prime})}\varphi_{s}(\bm{x}^{\prime})}{|\bm{r}-\bm{r}^{\prime}|}\,{\rm d}\bm{x}\,{\rm d}\bm{x}^{\prime}$
the two-electron integrals in this basis. Here, $\bm{x}=(\bm{r},\sigma)$ denotes the spatial and spin coordinates. Note that for the chosen basis, the complex conjugation in the above formula can be omitted as the basis functions are real. We also use the standard notation
$(pq||rs):=(pq|rs)-(ps|rq).$
The GHF energy functional $E^{\rm GHF}$ is given by
$E^{\rm GHF}(\gamma)={\rm Tr}(h\gamma)+\dfrac{1}{2}{\rm Tr}(G(\gamma)\gamma),$
where $G:{\mathbb{C}}_{\rm herm}^{N_{\rm b}\times N_{\rm b}}\to{\mathbb{C}}_{\rm herm}^{N_{\rm b}\times N_{\rm b}}$ is defined as $G(\gamma)=J(\gamma)-K(\gamma)$, with $J,K:{\mathbb{C}}_{\rm herm}^{N_{\rm b}\times N_{\rm b}}\to{\mathbb{C}}_{\rm herm}^{N_{\rm b}\times N_{\rm b}}$ the Coulomb and exchange functionals, respectively:
$[J(\gamma)]_{pq}:=\sum_{r,s=1}^{N_{\rm b}}(pq|rs)\gamma_{sr},\quad[K(\gamma)]_{pq}:=\sum_{\mu,\nu=1}^{N_{\rm b}}(ps|rq)\gamma_{sr}.$ (34)
It follows that $H_{\gamma}:=\nabla E^{\rm GHF}(\gamma)=h+G(\gamma)$ is the Fock matrix, that (27) coincides with the usual TDHF equations, and (28) with the McWeeny formulation of the HF equations. In GHF theory, the no-unfilled shell property proved in Ref. 54 in the continuous setting also holds in the discretized setting. We therefore have that in the chosen basis of occupied/virtual real molecular spin-orbitals
$\gamma_{\star}=\left(\begin{array}[]{cc}I_{N}&0\\ 0&0\end{array}\right),\quad H_{\gamma_{\star}}=\left(\begin{array}[]{cc}{\rm diag}(\varepsilon_{1},\ldots,\varepsilon_{N})&0\\ 0&{\rm diag}(\varepsilon_{N+1},\ldots,\varepsilon_{N_{\rm b}})\end{array}\right),$ (39)
$\mbox{with}\quad\varepsilon_{1}\leq\cdots\leq\varepsilon_{N}<\varepsilon_{N+1}\leq\cdots\leq\varepsilon_{N_{\rm b}}.$ (40)
Let us now show that the Kähler formalism provide a simple alternative to Casida’s derivation of LR excitation energies for mean-field models. Following the systematic strategy described in Section 2, it indeed suffices to select a basis
$\left(u_{1},\ldots,u_{N(N_{\rm b}-N)},J_{\gamma_{\star}}u_{1},\ldots,J_{\gamma_{\star}}u_{N(N_{\rm b}-N)}\right)$
of $T_{\gamma_{\star}}\mathcal{M}_{\mathbb{R}}$, build the matrix $\mathfrak{H}_{\star}$ defined in Eq. (11) with entries given by Eqs. (12)-(15), and compute the symplectic eigenvalues of $\mathfrak{H}_{\star}$. Choosing the canonical basis of $T_{\gamma_{\star}}\mathcal{M}_{\mathbb{R}}$, that is
$u_{[ia]}:=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&e_{[ia]}\\ e_{[ia]}^{T}&0\end{array}\right),\quad J_{\gamma_{\star}}u_{[ia]}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&ie_{[ia]}\\ -ie_{[ia]}^{T}&0\end{array}\right),\quad\mbox{with}\quad[e_{[ia]}]_{jb}=\delta_{(ia),(jb)},$
for $1\leq i\leq N$ and $N+1\leq j\leq N_{\rm b}$, we deduce from Eq. (30), the fact that $D^{2}E^{\rm GHF}(\gamma)(Q)=G(Q)$, and the real character of the spin-orbitals that
$\mathfrak{H}_{\star}^{\rm GHF}=\left(\begin{array}[]{cc}A+B&0\\ 0&A-B\end{array}\right),$ (41)
where the entries of the matrices $A$ and $B$ are given respectively by
$A_{ai,bj}:=(\varepsilon_{a}-\varepsilon_{i})\delta_{ij}\delta_{ab}+(ai||jb),\quad B_{ai,bj}:=(ai||bj).$
These matrices are the GHF analogues of the $A$ and $B$ matrices in LR-TDDFT Casida’s equations. The linearized dynamics Hamiltonian is therefore separable and the matrix $\widetilde{\Omega}$ in Eq. (18) then reads in this specific case
$\widetilde{\Omega}=\left((A+B)^{1/2}(A-B)(A+B)^{1/2}\right)^{1/2},$
an expression in agreement with Ref. 55 Eqs. (II.78)-(II.79).
4 Theoretical analysis in the weakly-interacting regime
In this part, we aim to compare theoretically, in the weakly-interacting regime, the CP and LR formalisms for computing the electronic excitation energies at the FCI and HF levels of theory. We consider a family of Hamiltonians parametrized by an additional coupling parameter $\lambda$ that modulates the degree of interaction within the energy functional. In the non-interacting limit ($\lambda=0$), HF and FCI lead to the same ground and excited state energies, so that CP-HF provides the same sequence of excitation energies as FCI, while LR-HF returns the subsequence of excitation energies corresponding to single excitations. Under a nondegeneracy assumption, we are able to compute the FCI (Section 4.1), as well as CP-HF and LR-HF (Section 4.2) excitation energies up to second-order errors in $\lambda$ in the weakly-interacting regime ($|\lambda|\ll 1$). To our knowledge, this is the first study aiming at comparing the CP and LR methodologies from a theoretical point of view. These theoretical findings are substantiated by numerical results in Section 5. We shall remark that although the cases of $\lambda\neq 1$ are non-physical, our analysis could offer potential insights into the study of more complicated real molecules.
In this section, we consider a system with $N=2N_{\rm p}$ electrons ($N_{\rm p}$ is the number of electron pairs) in $N_{\rm b}$ atomic orbitals (hence $2N_{\rm b}$ atomic spin-orbitals) and the family of second-quantized Hamiltonians on the Fock space ${\rm Fock}({\mathbb{C}}^{2N_{\rm b}})$ defined by
$\widehat{H}_{\lambda}:=\widehat{h}+\lambda\widehat{V}=\sum_{\kappa,\lambda=1}^{N_{\rm b}}h_{\kappa\lambda}\sum_{\sigma\in\{\alpha,\beta\}}\widehat{a}_{\kappa\sigma}^{\dagger}\widehat{a}_{\lambda\sigma}+\frac{\lambda}{2}\sum_{\kappa,\lambda,\mu,\nu=1}^{N_{\rm b}}\sum_{\sigma,\sigma^{\prime}\in\{\alpha,\beta\}}(\kappa\sigma,\mu\sigma|\lambda\sigma^{\prime},\nu\sigma^{\prime})\widehat{a}_{\kappa\sigma}^{\dagger}\widehat{a}_{\lambda\sigma^{\prime}}^{\dagger}\widehat{a}_{\nu\sigma^{\prime}}\widehat{a}_{\mu\sigma},$
where $\lambda\geq 0$ is a coupling parameter. Let $(\phi_{n}^{0})_{1\leq n\leq N_{\rm b}}$ be an orthonormal basis of ${\mathbb{C}}^{N_{\rm b}}$ diagonalizing the one-electron Hermitian matrix $h$, such that
$h\phi_{n}^{0}=\varepsilon_{n}^{0}\phi_{n}^{0}\quad\text{with}\quad\varepsilon_{1}^{0}\leq\cdots\leq\varepsilon_{N_{\rm b}}^{0}.$
To avoid technicalities, we assume that
the eigenvalues of $h$ are nondegenerate and the $\binom{N_{\rm b}}{N_{\rm p}}$ real numbers obtained by summing up $N_{\rm p}$ of the $N_{\rm b}$ eigenvalues of $h$ are all different.
4.1 Analysis at the FCI level of theory
It follows from Assumption (H) that the only degeneracies in the eigenstates of $\widehat{H}_{0}$ in the $N$-particle sector are due to degenerate spin states. Restricting our analysis to the $S_{z}=0$ spin sector, each state has $N_{\rm p}$ spin-up electrons and $N_{\rm p}$ spin-down electrons, the $N$-electron ground state of $\widehat{H}_{0}$ is a nondegenerate singlet state given by
$\Phi_{0}^{(0)}:=\frac{1}{\sqrt{N!}}\;(\phi_{1}^{0}\otimes\alpha)\wedge(\phi_{1}^{0}\otimes\beta)\wedge\cdots\wedge(\phi_{N_{\rm p}}^{0}\otimes\alpha)\wedge(\phi_{N_{\rm p}}^{0}\otimes\beta),$
and the first excited state is two-fold degenerate, spanned by
$\Phi^{(0)}_{\alpha}$ $:=\frac{1}{\sqrt{N!}}\;(\phi_{1}^{0}\otimes\alpha)\wedge(\phi_{1}^{0}\otimes\beta)\wedge\cdots\wedge(\phi_{N_{\rm p}-1}^{0}\otimes\alpha)\wedge(\phi_{N_{\rm p}-1}^{0}\otimes\beta)\wedge(\phi_{N_{\rm p}+1}^{0}\otimes\alpha)\wedge(\phi_{N_{\rm p}}^{0}\otimes\beta),$
$\Phi^{(0)}_{\beta}$ $:=\frac{1}{\sqrt{N!}}\;(\phi_{1}^{0}\otimes\alpha)\wedge(\phi_{1}^{0}\otimes\beta)\wedge\cdots\wedge(\phi_{N_{\rm p}-1}^{0}\otimes\alpha)\wedge(\phi_{N_{\rm p}-1}^{0}\otimes\beta)\wedge(\phi_{N_{\rm p}}^{0}\otimes\alpha)\wedge(\phi_{N_{\rm p}+1}^{0}\otimes\beta).$
Note that $\Phi^{(0)}_{\alpha}$ and $\Phi^{(0)}_{\beta}$ are not eigenfunctions of $\bm{S}^{2}$, but that $\Psi_{\rm S}^{(0)}:=(\Phi^{(0)}_{\alpha}-\Phi^{(0)}_{\beta})/\sqrt{2}$ is a singlet state (the first-excited state in the $\bm{S}^{2}=0$ sector) and $\Psi_{\rm T}^{(0)}:=(\Phi^{(0)}_{\alpha}+\Phi^{(0)}_{\beta})/\sqrt{2}$ the $S_{z}=0$ component of a triplet state (the ground state in the $\bm{S}^{2}=2$ sector). The energies of the ground state $\Phi^{(0)}_{0}$ and the two-fold degenerate first excited state are respectively
$E^{(0)}_{0}=2\sum_{i=1}^{N_{\rm p}}\varepsilon_{i}^{0}\quad\mbox{and}\quad E^{(0)}_{\alpha}=E^{(0)}_{\beta}=E^{(0)}_{\rm S}=E^{(0)}_{\rm T}=E^{(0)}_{0}+\varepsilon_{N_{\rm p}+1}-\varepsilon_{N_{\rm p}}.$
It follows from standard nondegenerate perturbation theory that for small $\lambda$, the ground state of $\widehat{H}_{\lambda}$ is nondegenerate and its energy $E_{0}(\lambda)$ is a real-analytic function of $\lambda$. Likewise the first excited state in the $\bm{S}^{2}=0$ sector as well as the ground state in the $\bm{S}^{2}=2$ sector are nondegenerate, and their energies $E_{\rm S}(\lambda)$ and $E_{\rm T}(\lambda)$ are real-analytic functions of $\lambda$ for $|\lambda|\ll 1$. Consequently, so are the FCI excitation energies $\omega_{\rm S}^{\rm FCI}(\lambda):=E_{\rm S}(\lambda)-E_{0}(\lambda)$ and $\omega_{\rm T}^{\rm FCI}(\lambda):=E_{\rm T}(\lambda)-E_{0}(\lambda)$:
$\omega_{\rm S}^{\rm FCI}(\lambda)=\sum_{k=0}^{+\infty}\lambda^{k}\omega^{{\rm FCI},(k)}_{\rm S},\quad\omega_{\rm T}^{\rm FCI}(\lambda)=\sum_{k=0}^{+\infty}\lambda^{k}\omega^{{\rm FCI}(k)}_{\rm T},$
where
$\omega^{{\rm FCI},(0)}_{\rm S}=\omega^{{\rm FCI},(0)}_{\rm T}=\omega^{(0)}_{\rm HOMO-LUMO}:=\varepsilon_{N_{\rm p}+1}-\varepsilon_{N_{\rm p}}.$
The coefficients $\omega^{(1)}_{\rm S}$ and $\omega^{(1)}_{\rm T}$ can be easily obtained by first-order perturbation theory:
$\omega^{{\rm FCI},(1)}_{\rm S/T}=$ $\sum_{j=1}^{N_{\rm p}-1}((N_{\rm p}+1)\beta,(N_{\rm p}+1)\beta||j\beta,j\beta)-\sum_{j=1}^{N_{\rm p}-1}(N_{\rm p}\beta,N_{\rm p}\beta||j\beta,j\beta)$
$+\sum_{j=1}^{N_{\rm p}}((N_{\rm p}+1)\beta,(N_{\rm p}+1)\beta|j\alpha,j\alpha)-\sum_{j=1}^{N_{\rm p}}(N_{\rm p}\beta,N_{\rm p}\beta|j\alpha,j\alpha)$
±$\pm(N_{\rm p}\beta,(N_{\rm p}+1)\beta|(N_{\rm p}+1)\alpha,N_{\rm p}\alpha).$ (42)
4.2 Analysis at the UHF level of theory
Let us now consider the UHF approximation, for which the underlying manifold is $\mathcal{M}={\rm Gr}_{\mathbb{C}}(N_{\rm p},N_{\rm b})\times{\rm Gr}_{\mathbb{C}}(N_{\rm p},N_{\rm b})$. The UHF approximation of $\widehat{H}_{\lambda}$ gives rise to a parametrized family of energy functionals of the form
$E^{\rm UHF}(\lambda,\gamma_{\alpha},\gamma_{\beta})=$ ${\rm Tr\,}(h\gamma_{\alpha})+{\rm Tr\,}(h\gamma_{\beta})+\frac{\lambda}{2}{\rm Tr\,}(J(\gamma_{\alpha}+\gamma_{\beta})(\gamma_{\alpha}+\gamma_{\beta}))$
$\quad-\frac{\lambda}{2}{\rm Tr\,}(K(\gamma_{\alpha})\gamma_{\alpha})-\frac{\lambda}{2}{\rm Tr\,}(K(\gamma_{\beta})\gamma_{\beta}),$ (43)
where $J$ and $K$ are defined in Eq. (34).
CP-UHF. The critical points of $\mathcal{M}\ni(\gamma_{\alpha},\gamma_{\beta})\mapsto\mathcal{E}^{\rm UHF}(0,\gamma_{\alpha},\gamma_{\beta})={\rm Tr\,}(h\gamma_{\alpha})+{\rm Tr\,}(h\gamma_{\beta})\in{\mathbb{R}}$ are the $(\gamma_{\alpha\star},\gamma_{\beta\star})\in\mathcal{M}$ satisfying $[h,\gamma_{\alpha\star}]=[h,\gamma_{\beta\star}]=0$. Under the nondegeneracy assumption (H), these critical points are isolated and there are exactly $\binom{N_{\rm b}}{N_{\rm p}}\times\binom{N_{\rm b}}{N_{\rm p}}$ of them. The ground state is
$\left(\gamma_{[\emptyset,\emptyset]}^{(0)},\gamma_{[\emptyset,\emptyset]}^{(0)}\right)\quad\mbox{with}\quad\gamma_{[\emptyset,\emptyset]}^{(0)}=\sum_{j=1}^{N_{\rm p}}\phi_{j}^{0}{\phi_{j}^{0}}^{T},$ (44)
and the excited states are
$\left(\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}^{(0)},\gamma_{\beta,[I_{\beta},A_{\beta}]}^{(0)}\right)\quad\mbox{where}\quad\gamma_{\sigma,[I_{\sigma},A_{\sigma}]}:=\gamma_{[\emptyset,\emptyset]}^{0}-\sum_{i\in I_{\sigma}}\phi_{i}^{0}{\phi_{i}^{0}}^{T}+\sum_{a\in A_{\sigma}}\phi_{a}^{0}{\phi_{a}^{0}}^{T},$ (45)
for $I_{\sigma}\subset\{1,\ldots,N_{\rm p}\}$, $A_{\sigma}\subset\{N_{\rm p}+1,\ldots,N_{\rm b}\}$, $|I|=|A|\geq 1$. It follows from a simple application of the analytic implicit function theorem that under the nondegeneracy assumption (H), there exist $\lambda_{\rm c}>0$ and $\binom{N_{\rm b}}{N_{\rm p}}\times\binom{N_{\rm b}}{N_{\rm p}}$ real-analytic functions
$\left(\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}(\bullet),\gamma_{\beta,[I_{\beta},A_{\beta}]}(\bullet)\right):(-\lambda_{\rm c},\lambda_{\rm c})\to\mathcal{M},$
$I_{\alpha},I_{\beta}\subset\{1,\ldots,N_{\rm p}\}$, $A_{\alpha},A_{\beta}\subset\{N_{\rm p}+1,\ldots,N_{\rm b}\}$, $|I_{\alpha}|=|A_{\alpha}|$, $|I_{\beta}|=|A_{\beta}|$, such that for any $\lambda\in(-\lambda_{\rm c},\lambda_{\rm c})$, the function $\mathcal{M}\ni(\gamma_{\alpha},\gamma_{\beta})\mapsto\mathcal{E}(\lambda,\gamma_{\alpha},\gamma_{\beta})\in{\mathbb{R}}$ has exactly $\binom{N_{\rm b}}{N_{\rm p}}\times\binom{N_{\rm b}}{N_{\rm p}}$ critical points, all isolated, nondegenerate, and given by $\left(\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}(\lambda),\gamma_{\beta,[I_{\beta},A_{\beta}]}(\lambda)\right)$. Note that, by symmetry, $\gamma_{\alpha,[\emptyset,\emptyset]}(\lambda)=\gamma_{\beta,[\emptyset,\emptyset]}(\lambda)=:\gamma_{[\emptyset,\emptyset]}(\lambda)$, i.e., the UHF ground state coincides with the RHF ground state.
For $\lambda\in(-\lambda_{\rm c},\lambda_{\rm c})$, the excitation energies obtained by CP-UHF are thus the $\binom{N_{\rm b}}{N_{\rm p}}\times\binom{N_{\rm b}}{N_{\rm p}}-1$ positive numbers
$\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF}}(\lambda):=E^{\rm UHF}\left(\lambda,\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}(\lambda),\gamma_{\beta,[I_{\beta},A_{\beta}]}(\lambda)\right)-E^{\rm UHF}\left(\lambda,\gamma_{[\emptyset,\emptyset]}(\lambda),\gamma_{[\emptyset,\emptyset]}(\lambda)\right)$ (46)
with $I_{\sigma}\subset\{1,\ldots,N_{\rm p}\}$, $A_{\sigma}\subset\{N_{\rm p}+1,\ldots,N_{\rm b}\}$, $|I_{\sigma}|=|A_{\sigma}|$, $|I_{\alpha}|+|I_{\beta}|=|A_{\alpha}|+|A_{\beta}|\geq 1$. In addition, the functions $\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF}}:(-\lambda_{\rm c},\lambda_{\rm c})\to{\mathbb{R}}$ are real-analytic and it holds
$\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF}}(\lambda)=\sum_{k=0}^{+\infty}\lambda^{k}\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF},(k)},$
with
$\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF},(0)}$ $=\sum_{a\in A_{\alpha}\cup A_{\beta}}\varepsilon_{a}^{0}-\sum_{i\in I_{\alpha}\cup I_{\beta}}\varepsilon_{i}^{0},$
$\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF},(1)}$ $=\frac{\partial E^{\rm UHF}}{\partial\lambda}\left(0,\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}^{(0)},\gamma_{\beta,[I_{\beta},A_{\beta}]}^{(0)}\right)-\frac{\partial E^{\rm UHF}}{\partial\lambda}\left(0,\gamma^{(0)}_{[\emptyset,\emptyset]},\gamma^{(0)}_{[\emptyset,\emptyset]}\right).$
The first-order variation of the excitation energy for the first excited state, given by the HOMO-LUMO excitation is
$\omega^{\text{CP-UHF},(1)}_{\rm HOMO-LUMO}:=$ $\omega^{\text{CP-UHF},(1)}_{[\emptyset,\emptyset],[N_{\rm p},N_{\rm p}+1]}=\omega^{\text{CP-UHF},(1)}_{[N_{\rm p},N_{\rm p}+1],[\emptyset,\emptyset]}$
$=$ $\sum_{i=1}^{N_{\rm p}-1}((N_{\rm p}+1)\sigma,(N_{\rm p}+1)\sigma||i\sigma,i\sigma)+\sum_{i=1}^{N_{\rm p}}((N_{\rm p}+1)\beta,(N_{\rm p}+1)\beta|i\alpha,i\alpha)$
$-\sum_{i=1}^{N_{\rm p}-1}(N_{\rm p}\sigma,N_{\rm p}\sigma||i\sigma,i\sigma)-\sum_{i=1}^{N_{\rm p}}(N_{\rm p}\beta,N_{\rm p}\beta|i\alpha,i\alpha)$ (47)
for $\sigma=\alpha$ or $\beta$.
LR-UHF. Let us now turn to the LR-UHF approximation of excitation energies. First, as the complex dimension of the manifold $\mathcal{M}$ is $2N_{\rm p}(N_{\rm b}-N_{\rm p})$, there are at most $2N_{\rm p}(N_{\rm b}-N_{\rm p})$ of them. For $\lambda=0$, they can be interpreted as single excitations:
$\omega_{[i,a]}^{{\rm LR-UHF},(0)}:=\omega_{[i,a],[\emptyset,\emptyset]}^{{\rm LR-UHF},(0)}=\omega_{[\emptyset,\emptyset],[i,a]}^{{\rm LR-UHF},(0)}=\varepsilon_{a}^{0}-\varepsilon_{i}^{0},\quad 1\leq i\leq N_{\rm p},\quad N_{\rm p}+1\leq a\leq N_{\rm b}.$
Using again the analytic implicit functional theorem, it can be shown that for $\lambda$ small enough, the excitation energies computed by linear response are given by real-analytic functions $(-\lambda_{\rm c},\lambda_{\rm c})\ni\lambda\mapsto\omega_{[i,a]}^{{\rm LR-UHF}}(\lambda)\in{\mathbb{R}}$ such that $\omega_{[i,a]}^{{\rm LR-UHF}}(0)=\omega_{[i,a]}^{{\rm LR-UHF},(0)}$. In particular, the first excited state corresponds to the HOMO-LUMO transition and the first-order variation is given by
$\omega^{\rm LR-UHF,(1)}_{\rm HOMO-LUMO}:=\omega^{\rm LR-UHF,(1)}_{[N_{\rm p},N_{\rm p}+1]}$
$=$ $\sum_{j=1}^{N_{\rm p}-1}((N_{\rm p}+1)\beta,(N_{\rm p}+1)\beta||j\beta,j\beta)-\sum_{j=1}^{N_{\rm p}-1}(N_{\rm p}\beta,N_{\rm p}\beta||j\beta,j\beta)$
$+\sum_{j=1}^{N_{\rm p}}((N_{\rm p}+1)\beta,(N_{\rm p}+1)\beta|j\alpha,j\alpha)-\sum_{j=1}^{N_{\rm p}}(N_{\rm p}\beta,N_{\rm p}\beta|j\alpha,j\alpha)$
$-(N_{\rm p}\beta,(N_{\rm p}+1)\beta|(N_{\rm p}+1)\alpha,N_{\rm p}\alpha).$ (48)
Notice that this expression coincides with the one found in the FCI case, in Eq. (42). Therefore, in the non-interacting limit, FCI and UHF-LR are equivalent in computing single excitation energies, while they differ in the interacting case.
5 Numerical simulations
We perform proof-of-concept numerical simulations to support our theoretical findings. In Section 5.1, we compare the CP and LR formalisms in UHF theory for the H2, H4, and H2O molecules, with a varying coupling parameter $\lambda$. The exact FCI theory is taken as reference. In particular, we check the analytical first-order derivatives of the excitation energies at $\lambda=0$ given by Eqs. (47) and (48) vs the numerical ones obtained by local linear fits. In Section 5.2, we demonstrate that multiple critical points can arise even for the H4 molecule (rectangular geometry) at the UHF/STO-3G level of theory. The energies of the critical points and their overlaps with the exact FCI excited states are shown as functions of $\lambda$. We note that in this section, only the index-1 saddle points (namely, the critical points where the Riemannian Hessian has exactly one negative eigenvalue) are considered in the CP formalism and they are computed numerically by using a constrained gentlest ascent dynamics, which is described in detail in another ongoing work. The approximate excitation energies in the LR formalism are found by solving Casida’s equations 23, 24, 25 using the open-source package PySCF 56.
5.1 Comparison of CP and LR formalisms within the UHF theory
In this Section, numerical experiments performed in the UHF framework are shown. Calculations have been performed with the minimal basis set STO-3G 57 and with the 3-21G basis 58, for the systems H2, H4 in a linear and a rectangular geometry, and H2O, with $\lambda$ in the range $[0,1]$. For H2, a bond length of 1 a.u. was considered; for the H4 molecule, the linear geometry considers a bond length of 0.875 Å (Ref. 59); for the rectangular geometry, sides of 1 Å and 2 Å are considered. For the water molecule, the coordinates [O (0.00000000 0.00000000 0.11993333) a.u.], [H ($00$1.43497461 -0.95171452) a.u.], [H (0.00000000 1.43497461 -0.95171452) a.u.] were taken. We compare the first-order excitation energies of the HOMO-LUMO transition, obtained both with the LR and CP approaches. We note that for the CP results in this subsection, only the lowest excited states at $\lambda=0$, which correspond to the HOMO-LUMO excitations, have been tracked across the range of $\lambda$. When $\lambda$ increases, multiple critical points (with the same index) may emerge; see the discussions in Section 5.2.
Table 1: First-order excitation energies for the HOMO-LUMO excitations of the molecules H2, H4 in linear and rectangular geometry, and H2O, obtained in the non-interacting regime of $\lambda\xrightarrow[]{}0$. For both CP and LR, the results obtained by the analytic expressions Eqs. (47) and (48), denoted here by $\omega_{\rm anal}^{{\rm CP},(1)}$ and $\omega_{\rm anal}^{{\rm LR},(1)}$, and numerically using a five-point scheme linear fit, denoted here by $\omega_{\rm num}^{{\rm CP},(1)}$ and $\omega_{\rm num}^{{\rm LR},(1)}$, are listed. The difference $\Delta\omega_{\rm anal}^{(1)}=\omega_{\rm anal}^{{\rm CP},(1)}-\omega_{\rm anal}^{{\rm LR},(1)}$ between the two approaches is also shown.
Molecule (Basis) $\omega_{\rm num}^{{\rm CP},(1)}$ $\omega_{\rm anal}^{{\rm CP},(1)}$ $\omega_{\rm num}^{{\rm LR},(1)}$ $\omega_{\rm anal}^{{\rm LR},(1)}$ $\Delta\omega_{\rm anal}^{(1)}$
H2 (STO-3G) -0.012580 -0.012586 -0.182839 -0.182827 0.170241
H2 (3-21G) -0.330913 -0.330929 -0.422845 -0.422889 0.091960
H4 lin. (STO-3G) -0.107158 -0.107347 -0.245692 -0.245812 0.138465
H4 lin. (3-21G) -0.207375 -0.207529 -0.332364 -0.332485 0.124956
H4 rect. (STO-3G) 0.052118 0.052118 -0.040455 -0.040246 0.092364
H4 rect.(3-21G) -0.061798 -0.061775 -0.133340 -0.133204 0.071429
H2O (STO-3G) -3.494042 -3.494149 -3.497100 -3.497208 0.003059
The results displayed in Table 1 show for all examples a good agreement between the analytical derivatives obtained in Section 4 and linear fits of these quantities using the five points $(j\cdot\delta\lambda)_{0\leq j\leq 4}$, with $\delta\lambda=10^{-4}$ chosen small enough to be compatible with the limit $\lambda\xrightarrow[]{}0$, but not too small to avoid numerical instabilities. Furthermore, we can observe that, for the examples considered here, the difference calculated as $\omega_{\mathrm{anal}}^{{\rm CP},(1)}-\omega_{\mathrm{anal}}^{{\rm LR},(1)}$ is always positive, showing that the LR result is always smaller than the CP one for the explored systems.
Figure 2: Dependence on the $\lambda$-parameter of the excitation energy, calculated on the H2 molecule, with the 3-21G basis set.
Figure 3: Dependence on the $\lambda$-parameter of the excitation energy, calculated on the H4 molecule, in the linear (a) and rectangular (b) geometries, with the 3-21G basis set.
The two approaches CP and LR, at the UHF level of theory can be compared with the exact FCI results. The results are shown in Figs. 2-4, for the molecules H2 and H4 in the linear and rectangular geometry, with the 3-21G basis set 58.
Figure 4: Dependence on the $\lambda$-parameter of the excitation energy, calculated on the H2O, with the STO-3G basis set. In Fig. 4(b), a zoom on the differences between the different approaches in the region around $\lambda=1$ is plotted.
Figure 5: Dependence on the $\lambda$-parameter of the excitation energy, calculated on H2O. In Fig. 5(a) the 6-31G basis set is used, while in Fig. 5(b), the cc-pVDZ basis set is used. At $\lambda=0.2$ and $\lambda=0.6$, the TD-HF algorithm in PySCF was not able to converge due to high linear dependencies.
For all cases, the LR results are closer to the FCI ones. The plots also show that, in the non-interacting limit ($\lambda=0$), the LR formalism coincides with FCI as expected from Eq. (48), while the two approaches differ in the interacting case. In the water example with the STO-3G (Fig. 4) 57, 6-31G 60, 61, and cc-pVDZ 62 basis sets (Fig. 5), the three methods show small discrepancies across the range until $\lambda$ approaches 1.
5.2 Multiple critical points within the UHF theory for H4
Due to the nonlinearities induced by the UHF approximation, there can be multiple critical points (in particular, index-1 saddle points) over the UHF electronic landscape. In this subsection, we consider the H4 molecule in a rectangular configuration with side lengths of 1 Å and 1.1 Å, at the UHF/STO-3G 57 level of theory. The Hamiltonian is parametrized by the coupling parameter $\lambda$, as in Section 4.1. For each $\lambda\in[0,2]$ with a spacing of $0.05$, we find the index-1 saddle points by running the saddle search algorithm from 50 different randomly generated initial points over the Grassmannian.
Fig. 6 shows the energies of the index-1 saddle points which appear in the range $\lambda\in[0,2]$.
Figure 6: Dependence on the $\lambda$-parameter of the energy of the index-1 saddle points, calculated on H4 in a rectangular configuration with sides of 1 Å and 1.1 Å. The UHF/STO-3G level of theory has been used. Calculations have been made in steps of $\Delta\lambda=0.2$.
We see that in this range, at most three critical points are found. In the range $\lambda\in[0,0.35]$, only one index-1 saddle point is found, in $\lambda\in[0.40,0.60]$ two index-1 saddle points are found, and from there with increasing $\lambda$, three index-1 saddle points are observed. At $\lambda\approx 1.68$ the lines corresponding to the second and third saddle points are crossing, showing that the ordering in the energies of the states might not be preserved.
It is known that not all index-1 saddle points provide suitable approximations for excited states. An interpretation of the critical points can be gained through projection onto the FCI states, as these form a complete basis. Our calculations consider the case of $S_{z}=0$, therefore fixing $N_{\alpha}=N_{\beta}=2$. In the present case of a minimal basis set, a total number of 36 FCI states can be built. The projection of the wave functions corresponding to the critical points can be written as
|$\ket{\Psi^{\mathrm{CP-HF}}(\lambda)}=\sum_{I}c_{I}(\lambda)\ket{\Psi_{I}^{\mathrm{FCI}}(\lambda)}.$
The magnitudes of the $c_{I}(\lambda)$ coefficients indicate the extent to which the critical points correspond to the $I$-th FCI state. For the sake of clarity, we will include only the eight lowest-lying FCI states, as the contributions of the higher-energy states are relatively small. We include the evolution of the considered FCI excited energies with the coupling parameter in Fig. 7(a).
Figure 7: (a) Dependence on the $\lambda$-parameter of the FCI energies, and (b-d) dependence on the $\lambda$-parameter of the overlaps of the UHF critical points with the FCI states, calculated on H4 in a rectangular configuration with sides of 1 Å and 1.1 Å. Calculations have been made in steps of $\Delta\lambda=0.05$.
We label these states as $\Psi_{0}^{\mathrm{FCI}}(\lambda)$, $\Psi_{1}^{\mathrm{FCI}}(\lambda)$, $\ldots$, $\Psi_{7}^{\mathrm{FCI}}(\lambda)$ in accordance to their energy ordering in the non-interacting case. At $\lambda=0$, we observe that the pairs ($\Psi_{1}^{\mathrm{FCI}}(0)$, $\Psi_{2}^{\mathrm{FCI}}(0)$), ($\Psi_{4}^{\mathrm{FCI}}(0)$, $\Psi_{5}^{\mathrm{FCI}}(0)$), and ($\Psi_{6}^{\mathrm{FCI}}(0)$, $\Psi_{7}^{\mathrm{FCI}}(0)$) are degenerate. These degeneracies can be understood by considering the occupation of the molecular orbitals. In the non-interacting case, the lowest eight FCI states are either singlets or triplets and are therefore either single Slater determinants or linear combinations of two Slater determinants:
|$\ket{\Psi_{0}^{\rm FCI}(0)}$ =|$=\ket{\phi_{1}^{\alpha}\phi_{1}^{\beta}\phi_{2}^{\alpha}\phi_{2}^{\beta}}$
(|$(\ket{\Psi_{1}^{\rm FCI}(0)},\ket{\Psi_{2}^{\rm FCI}(0)})$ $=\frac{1}{\sqrt{2}}(\ket{\phi_{1}^{\alpha}\phi_{1}^{\beta}\phi_{2}^{\alpha}\phi_{3}^{\beta}}\pm\ket{\phi_{1}^{\alpha}\phi_{1}^{\beta}\phi_{2}^{\beta}\phi_{3}^{\alpha}})$
|$\ket{\Psi_{3}^{\rm FCI}(0)}$ =|$=\ket{\phi_{1}^{\alpha}\phi_{1}^{\beta}\phi_{3}^{\alpha}\phi_{3}^{\beta}}$
(|$(\ket{\Psi_{4}^{\rm FCI}},\ket{\Psi_{5}^{\rm FCI}(0)})$ $=\frac{1}{\sqrt{2}}(\ket{\phi_{1}^{\alpha}\phi_{2}^{\beta}\phi_{2}^{\alpha}\phi_{3}^{\beta}}\pm\ket{\phi_{1}^{\beta}\phi_{2}^{\beta}\phi_{2}^{\alpha}\phi_{3}^{\alpha}})$
(|$(\ket{\Psi_{6}^{\rm FCI}(0)},\ket{\Psi_{7}^{\rm FCI}(0)})$ $=\frac{1}{\sqrt{2}}(\ket{\phi_{1}^{\alpha}\phi_{1}^{\beta}\phi_{2}^{\alpha}\phi_{4}^{\beta}}\pm\ket{\phi_{1}^{\alpha}\phi_{1}^{\beta}\phi_{2}^{\beta}\phi_{4}^{\alpha}})$
The degeneracies arise from the fact that at $\lambda=0$, the singlet state and the triplet state of $S_{z}=0$ have the same energy, as discussed in Section 4.1.
For $\lambda>0$, it can be observed that the degeneracy is lifted, as the singlet and triplet states no longer give the same energy. With increasing interaction parameter $\lambda$, the FCI states are observed to cross. In particular, $\Psi_{2}^{\mathrm{FCI}}$ (green line in Fig. 7(a)) crosses consecutively with the states $\Psi_{3}^{\mathrm{FCI}}$, $\Psi_{4}^{\mathrm{FCI}}$, $\Psi_{5}^{\mathrm{FCI}}$, $\Psi_{6}^{\mathrm{FCI}}$ (respectively, red, purple, brown, and pink lines in Fig. 7(a)) at $\lambda\approx 0.77$, 1.33, 1.49, and 1.94 respectively, as indicated by the dot, square, star, and diamond markers in the figure.
In Figs. 7(b)-7(d), the $c_{I}$ coefficients corresponding to the three index-1 saddle points are plotted. In Fig. 7(b), the analysis is performed for the lowest-lying critical point (CP1, blue line in Fig. 6). This saddle point appears only from $\lambda\approx 0.4$ and shows to be dominated by the FCI ground state. Therefore, this critical point can be interpreted as a spurious saddle point arising from the use of an approximate wave function method (UHF), which introduces nonlinearities. The existence of spurious critical points in nonlinear approximation models has been reported many times since the early days of quantum chemistry, and the study of this phenomenon is still an active field of research (cf. e.g., the recent work by Marie and Burton in Ref. 7 on CASSCF). Fig. 7(c) shows the same analysis for the second critical point (CP2, orange line in Fig. 6). This state remains a combination of $\Psi_{1}^{\rm FCI}$ and $\Psi_{2}^{\rm FCI}$. In particular, at $\lambda=0$, we have |$\ket{\Psi_{2}^{\mathrm{HF,CP}}(0)}=\frac{1}{\sqrt{2}}(\ket{\Psi_{1}^{\mathrm{FCI}}(0)}+\ket{\Psi_{2}^{\mathrm{FCI}}(0)})$ at $\lambda=0$. As the two states $\Psi_{1}^{\rm FCI}$ and $\Psi_{2}^{\rm FCI}$ are degenerate at $\lambda=0$ but belong to different spin-symmetry sectors, this critical point is a symmetry-broken solution, formed by a linear combination of the singlet first excited state $\Psi_{1}^{\rm FCI}$ and the triplet ground state $\Psi_{2}^{\rm FCI}$.
In Fig. 7(d), the overlap of the third critical point (CP3, green line in Fig. 6) is shown. It is observed that this critical point, which arises only from $\lambda=0.60$, has a large overlap with the state $\Psi_{3}^{\mathrm{FCI}}$. In particular, for $\lambda\in[0.60,1]$ the overlap is very close to one, but then the saddle point becomes a linear combination of states $\Psi_{3}^{\mathrm{FCI}}$ and $\Psi_{5}^{\mathrm{FCI}}$.
The analysis confirms once more that the interpretation of critical points as excited states should therefore be performed with care. On the reported example, only one critical point of a Morse index 1 (CP2 in Fig. 6) can be interpreted as a proper approximation of the first excited state in the UHF setting, while the other two are spurious states arising from the intrinsic nonlinearity of the UHF theory.
Acknowledgments
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement EMC2 No 810367). This work has benefited from French State support managed by ANR under the France 2030 program through the MaQui CNRS Risky and High-Impact Research programme (RI)2 (grant agreement ANR-24-RRII-0001). The authors are grateful to Théo Duez, Asbjörn Lauritsen, Tony Lelièvre, Mathieu Lewin, Filippo Lipparini, Tommaso Nottoli, Panos Parpas, Julien Toulouse, and Jari van Gog for useful discussions.